# ManiCM: Real-time 3D Diffusion Policy via Consistency Model for Robotic Manipulation

### v1

@@ 3.3 Manipulation Consistency Distillation
To train ManiCM, we adopt the method of consistency distillation, as depicted in Figure 2. Here $\boldsymbol{a}_{0}$ represents the sequence of robotic arm actions, which comprises multiple actions, with the dimensionality determined by the specific task. $p_{0}$ denotes the point cloud obtained from images and subjected to downsampling, while $q_{0}$ signifies the current pose of the robotic arm from the offline dataset. These inputs are processed through an encoder to obtain compact representations, which are then used as conditional inputs for the teacher network, online network, and target network. Given a ground-truth expert action $\boldsymbol{a}_{0}$, we first obtain the noisy action $\boldsymbol{a}_{n+k}$ by conducting a forward diffusion operation with $n+k$ steps to add noise on $\boldsymbol{a}_{0}$, where $k$ is the skipping interval mentioned in Section 3.1. Subsequently, the noisy action $\boldsymbol{a}_{n+k}$ is feed-forwarded to a frozen teacher network and the online network to decode the clean $\hat{\boldsymbol{a}}_{0}^{*}$ and $\hat{\boldsymbol{a}}_{0}$, respectively. In order to make sure the self-consistency property Equation 1, we design a target network cloned from the online network, and we expect the denoised outputs of the online network and the target network to be aligned. Specifically, the target network generates $\hat{\boldsymbol{a}}_{0}^{-}$ utilizing $\boldsymbol{a}_{n}$ obtained from processing $\hat{\boldsymbol{a}}_{0}^{*}$ with a $k$-step ODE solver $\bf\Phi$ (e.g., DDIM [36]), such that $\hat{\boldsymbol{a}}_{0}$ and $\hat{\boldsymbol{a}}_{0}^{-}$ are obtained by the consistency function $\textbf{{f}}(\cdot,\cdot)$ in  Equation 2. Therefore, the consistency distillation loss is expressed as,
@@
where ${\bf\Phi}(\cdot,\cdot)$ is the update funciton of a $k$-step ODE solver applied to PF-ODE. As outlined in Section 3.1, we utilize EMA to update the parameters of the target network from the parameters of the online network. We refer to the pre-trained manipulation 3D diffusion model, such as DP3 [48], as the "teacher network". In practice, both the online network and target network are initialized with the parameters of the teacher network.
@@
Compared to other manipulation diffusion models, our ManiCM model achieves the fastest runtime ($\sim$16ms per action) during the inference phase by sampling high-quality actions in just one step, which is illustrated in Table 1.
4 Experiments
In this section, we first introduce the experiment setup including datasets, baseline methods, evaluation metrics and implementation details (Section 4.1). Then we compare our method with the state-of-the-art approaches to show the superiority in runtime efficiency and success rate (Section 4.2). Subsequently, we conduct an ablation study to explore different design choices in our ManiCM (Section 4.3). We also illustrate the visualization results to depict our intuition (Section 4.4). Further results can be found in the supplementary material.
4.1 Experimental Setup
Datasets. We conduct our experiments in the well-recognized MetaWorld [45] and Adroit [31] benchmarks, resulting in a total of $31$ tasks. These tasks range from simple pick-and-place tasks to more challenging scenarios such as dexterous manipulation, which ensure that the model is effective across various scenarios. Specifically, MetaWorld tasks are grouped into different difficulty levels from easy to very hard according to [33]. For offline training datasets, we obtain expert demonstrations through heuristic policies in MetaWorld, and acquire them utilizing the reinforcement learning method VRL3 [38] in Adroit. We obtain a limited number of $10$ expert demonstrations for training in each benchmark, following [48] for fair comparisons.
Baselines. This work primarily emphasizes accelerating the inference speed of 3D modality in diffusion policy to achieve real-time robotic action generation. To this end, our main baseline is the state-of-the-art point cloud-based 3D diffusion policy (DP3 [48]). We reproduce DP3 with higher performance than the original paper, and denote it as DP3. Additionally, we compare our method with ‘Simple DP3’, a variation of DP3 with a lightweight policy backbone. Our baselines also include the well-known diffusion policy (DP [7]) for reference. To ensure fairness, the resolution of all images and depth maps remains consistent across all experiments.
Evaluation Metrics. Following [48], we assess $20$ episodes every $200$ training epochs and then calculate the average of the highest $5$ success rates, along with the average runtime per step. For each main experiment, we run $3$ random seeds with seed numbers 00, $1$, $2$ to mitigate the performance fluctuation. We also report the Number of Function Evaluations (NFE) as an additional device-agnostic metric for inference speed. For the success rate and runtime, the higher is the better. For the NFE, the lower the better.
Implementation Details. For ManiCM, an AdamW [19] optimizer with a batch size of 128 and a learning rate of 5e-5 is employed for 3K epochs training. We also adopt a cosine decay learning rate scheduler and 500 iterations of linear warm-up. The EMA rate is set to $\mu=0.95$ by default and we use DDIM-Solver [36] with a skipping step of $k=10$ as ODE Solvers. Unless otherwise specified, our experiments are conducted using $100$ training time steps. The teacher model employed in our experiments uses DDIM with an inference step of $10$. For the input, We utilize an observation steps value of $2$, indicating that we use the point clouds from the two most recent time steps as conditions. Similarly, with a horizon of $4$, we predict four consecutive actions during each iteration for long-horizon tasks, while only $2$ actions are executed per prediction. To focus on the interest area, the image size is cropped to $84\times 84$ pixels. The point cloud is downsampled to $512$ points using the Farthest Point Sampling (FPS) algorithm, while the dimension of agent poses and actions is determined by the specific task. Both point clouds and agent poses are encoded using an MLP encoder into embeddings of length $64$. For normalization, we limit the min and max of action and observation to $\left[-1,1\right]$ independently since the prediction of DDIM would clip the prediction to $\left[-1,1\right]$ to ensure training stability. We implement our model in PyTorch [28], and train the model on one NVIDIA RTX 4090 GPU. All compared models are evaluated on one NVIDIA RTX 4090 GPU to benchmark the decision runtime.
Table 1: Comparisons on Runtime. We evaluate $100$ episodes on $31$ challenging tasks from Adroit and Metaworld across $3$ random seeds and report the time consumption per step (s) with standard deviation. The second results are underlined and the best results are bold. ‘’ denotes the reproduced version. The performance of our ManiCM in one-step inference surpasses all state-of-the-art models, providing ample evidence for the effectiveness of consistency distillation.
DP [7] $10$ $162.6\scriptstyle{\pm 1.2}$
DP3 [48] $10$ $157.4\scriptstyle{\pm 1.5}$
DP3 $10$ $159.6$ $177.4$ $186.1$ $183.2$ $181.8$ $177.6\scriptstyle{\pm 1.4}$
Simple DP3 $10$ $95.5$ $123.5$ $119.2$ $124.6$ $124.7$ $120.6\scriptstyle{\pm 0.9}$
ManiCM ($4$-step) $4$ $56.3$ $66.8$ $68.2$ $67.9$ $69.4$ $\underline{66.2\scriptstyle{\pm 0.3}}$
Table 2: Comparisons on Success Rate. We evaluate $100$ episodes on $31$ challenging tasks from Adroit and Metaworld across $3$ random seeds and report the success rates (%) with standard deviation. The second results are underlined and the best results are bold. ‘’ denotes the reproduced version. The performance of our ManiCM in one-step inference surpasses all state-of-the-art models, providing ample evidence for the effectiveness of consistency distillation.
DP [7] $10$ $31.6$ $77.1$ $18.0$ $3.0$ $22.5$ $52.8\scriptstyle{\pm 3.6}$
DP3 [48] $10$ $68.3$ $86.0$ $25.0$ $13.3$ $44.0$ $65.9\scriptstyle{\pm 3.3}$
DP3 $10$ $76.1$ $91.7$ $46.3$ $44.8$ $62.3$ $77.5\scriptstyle{\pm 3.8}$
Simple DP3 $10$ $73.0$ $90.4$ $47.3$ $45.6$ $56.5$ $75.8\scriptstyle{\pm 3.3}$
ManiCM ($1$-step) $1$ $74.9$ $92.9$ $47.7$ $44.2$ $\mathbf{65.7}$ $\underline{78.5\scriptstyle{\pm 4.2}}$
4.2 Comparison with the State-of-the-Art Methods
Comparisons on Runtime. Table 1 illustrates the comparison of the runtime per step of each model. Our method achieves an impressive runtime around 16ms per step on average, which is $\sim$10$\times$ faster than the baseline method DP3. The acceleration comes from the manipulation consistency model, which can generate robot action in only one-step inference. Besides, the $4$-step inference version of ManiCM also improves the efficiency by $\sim$3$\times$, Moreover, the proposed ManiCM also surpasses the lightweight version Simple DP3 significantly, which shows that the speedup from reducing the sampling speed is a more efficient way than shrinking the model itself. The results demonstrate the superiority of our ManiCM in terms of runtime, which verifies our intuition that incorporating the consistency model boosts the inference speed in various robotic manipulation tasks.
Comparisons on Success Rate. Table 2 depicts the comparison of the success of each model. The results show that our method accelerates the diffusion process without any performance drop. Specifically, with only one-step inference, our ManiCM can approximate or even surpass the state-of-the-art model DP3 and Simple DP3 by $1.3$%. By distilling the diffusion process with consistency regularization, our method avoids overfitting to the training demonstrations. Especially, in the Metaworld Very Hard tasks, our method outperforms the state-of-the-art method with an absolute improvement of $5.5$%. Moreover, increasing the sampling steps in ManiCM yields even better performance on average ($79.0$% vs. $78.5$%). The above experimental results prove the effectiveness of handling robotic tasks with our manipulation consistency model.
@@ 4.2 Comparison with the State-of-the-Art Methods
Table 3: Ablation on Different Diffusion Steps of the Teacher Model. ‘SR’ refers to the success rate. $\rightarrow$’ means skipping from the original training denoising steps to $k$ steps in inference with DDIM-Solver.
@@
4.3 Ablation Studies
@@
Ablation on Different Diffusion Steps of the Teacher Model. We compare our ManiCM with the state-of-the-art method DP3 under different DDIM sampling steps in Table 3. We can conclude that: 1) The task success rates do not monotonically increase when taking more diffusion sampling steps in both DP3 and ManiCM, which indicates that most of the diffusion steps are redundant. As a result, we can distill the teacher model with multiple-step inference into an online model that generates high-quality action in only one sampling step. 2) For different diffusion sampling steps, our ManiCM can stably distill competitive capacity in task completion, where it accelerates the runtime efficiency by $\sim$57 $\times$ when the teacher is a DDIM with $50$ inference steps.
4.4 Qualitative Analysis
@@
5 Conclusion
In this paper, we have proposed a ManiCM agent that imposes the consistency constraint on the diffusion process, which brings 3D diffusion-based robotic manipulation to a real-time level. We design a manipulation consistency self-consistency function that predicts the action sample directly based on 3D point cloud condition. Subsequently, we implement a manipulation consistency distillation to ensure the action is denoised from any point along the ODE trajectory within only step inference, which avoids the iterative denoising procedure in the original diffusion model. Extensive experiments across various task suites verify the effectiveness and efficiency of our method.
6 Limitation
First, though MotionLCM accelerates diffusion-based manipulation policies with no performance drop in various robotic manipulation tasks, the performance bottleneck of ManiCM relies on the teacher model by consistency distillation. Second, the scalability of the consistency model to distill a large-scale pre-trained diffusion policy is still under-explored. We leave these issues for future work.
@@ Acknowledgement
Appendix A Appendix
A.1 Task Suite
We conduct experiments on a wide range of tasks to show the effectiveness of our method, including $3$ tasks from Adroit [31] and $28$ tasks from Metaworld [45]. These tasks varies in the robot model, action dimension and number of variations, which is summarized in Table 4.
Table 4: Task suite. We summarize the properties of the tasks we use in main experiments. ‘Dimension’ refers to the action dimension for the domain. ‘#Task’ is the number of tasks included in each benchmark. ‘#Demo’ means the number of expert demonstrations used for each task.
Task Suite ($31$ Tasks in total)
Domain Robot Model Dimension #Task #Demo
Adroit Shadow 28 3 10
MetaWorld Gripper 4 28 10
A.2 Addtional Implementation Details
For reproducibility, the training settings and hyperparameters used in ManiCM are shown in Table 5. Most of the hyperparameters are in line with previous works [7, 48] for fair comparisons.
Table 5: Hyperparameters.
Hyperparameter Value
Training Epoch $3$K
Image Resolution $84\times 84$
# of Points $512$
Batch Size $128$
Optimizer AdamW
Learning Rate $0.0005$
Historical Observations $2$
Horizon $4$
Training Timesteps of Teacher Model $100$
Inference Timesteps of Teacher Model $10$
EMA Rate $0.95$
A.3 Complete Results on $31$ Tasks
We provide the complete results on runtime and success rate compared with the state-of-the-art method DP3 [48] in Table 6 and Table 7 respectively, as a supplement to the main results in the paper.
Table 6: Complete Results on Runtime. We report the runtime taken per step for each task in detail. ‘ManiCM’ refers to the 1-step version of our method.
Adroit [31] Meta-World (Easy) [45]
Method/Task Hammer Door Pen Button Press Button Press Wall Coffee Button Dial Turn Door Close
DP3 [48] $192.4\scriptstyle{\pm 1.8}$ $107.3\scriptstyle{\pm 0}$ $179.0\scriptstyle{\pm 0.1}$ $175.3\scriptstyle{\pm 1.4}$ $187.1\scriptstyle{\pm 1.0}$ $187.7\scriptstyle{\pm 1.1}$ $155.2\scriptstyle{\pm 4.0}$ $186.3\scriptstyle{\pm 0.9}$
ManiCM $\mathbf{18.6\scriptstyle{\pm 0.2}}$ $\mathbf{15.6\scriptstyle{\pm 0.3}}$ $\mathbf{16.5\scriptstyle{\pm 0.0}}$ $\mathbf{18.3\scriptstyle{\pm 0.0}}$ $\mathbf{18.9\scriptstyle{\pm 0.0}}$ $\mathbf{19.4\scriptstyle{\pm 0.0}}$ $\mathbf{12.9\scriptstyle{\pm 0.2}}$ $\mathbf{18.9\scriptstyle{\pm 0}}$
Meta-World (Easy)
Method/Task Door Open Drawer Close Faucet Open Handle Pull Handle Press Side Handle Pull Side Lever Pull Plate Slide
DP3 [48] $157.6\scriptstyle{\pm 4.4}$ $128.7\scriptstyle{\pm 3.7}$ $171.6\scriptstyle{\pm 1.2}$ $180.9\scriptstyle{\pm 1.2}$ $183.5\scriptstyle{\pm 1.3}$ $182.8\scriptstyle{\pm 2.3}$ $180.9\scriptstyle{\pm 1.1}$ $191.6\scriptstyle{\pm 0.8}$
ManiCM $\mathbf{16.5\scriptstyle{\pm 0.3}}$ $\mathbf{15.8\scriptstyle{\pm 0.3}}$ $\mathbf{16.7\scriptstyle{\pm 0.2}}$ $\mathbf{18.1\scriptstyle{\pm 0.1}}$ $\mathbf{18.1\scriptstyle{\pm 0.1}}$ $\mathbf{18.1\scriptstyle{\pm 0.1}}$ $\mathbf{16.4\scriptstyle{\pm 0.2}}$ $\mathbf{19.9\scriptstyle{\pm 0.1}}$
Meta-World (Easy) Meta-World (Medium)
Method/Task Plate Slide Back Side Reach Reach Wall Window Close Peg Unplug Side Box Close Soccer Sweep Into
DP3 [48] $191.6\scriptstyle{\pm 0.4}$ $180.7\scriptstyle{\pm 1.1}$ $186.6\scriptstyle{\pm 0.4}$ $181.8\scriptstyle{\pm 2.1}$ $183.1\scriptstyle{\pm 1.5}$ $184.2\scriptstyle{\pm 1.5}$ $188.3\scriptstyle{\pm 1.3}$ $185.9\scriptstyle{\pm 0.4}$
ManiCM $\mathbf{19.1\scriptstyle{\pm 0.1}}$ $\mathbf{17.4\scriptstyle{\pm 0.2}}$ $\mathbf{16.3\scriptstyle{\pm 0.2}}$ $\mathbf{18.6\scriptstyle{\pm 0.1}}$ $\mathbf{17.0\scriptstyle{\pm 0.3}}$ $\mathbf{16.8\scriptstyle{\pm 0.3}}$ $\mathbf{19.4\scriptstyle{\pm 0.2}}$ $\mathbf{15.5\scriptstyle{\pm 0.1}}$
Meta-World (Hard) Meta-World (Very Hard)
Method/Task Hand Insert Pick Out of Hole Pick Place Shelf Place Stick Pull Stick Push Pick Place Wall
DP3 [48] $186.8\scriptstyle{\pm 0.8}$ $178.0\scriptstyle{\pm 0.9}$ $184.7\scriptstyle{\pm 1.2}$ $187.4\scriptstyle{\pm 0.4}$ $182.8\scriptstyle{\pm 1.6}$ $175.8\scriptstyle{\pm 1.8}$ $181.3\scriptstyle{\pm 2,1}$
ManiCM $\mathbf{16.3\scriptstyle{\pm 0.3}}$ $\mathbf{17.1\scriptstyle{\pm 0.2}}$ $\mathbf{17.6\scriptstyle{\pm 0.1}}$ $\mathbf{16.5\scriptstyle{\pm 0.3}}$ $\mathbf{17.7\scriptstyle{\pm 0.1}}$ $\mathbf{16.0\scriptstyle{\pm 0.1}}$ $\mathbf{16.0\scriptstyle{\pm 0.1}}$
Table 7: Complete Results on Success Rate. We report the success rate for each task in detail. ‘ManiCM’ refers to the 1-step version of our method.
Adroit [31] Meta-World (Easy) [45]
Method/Task Hammer Door Pen Button Press Button Press Wall Coffee Button Dial Turn Door Close
DP3 [48] $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{75\scriptstyle{\pm 6}}$ $\mathbf{54\scriptstyle{\pm 4}}$ $99\scriptstyle{\pm 1}$ $\mathbf{100\scriptstyle{\pm 0}}$ $91\scriptstyle{\pm 3}$ $80\scriptstyle{\pm 6}$ $\mathbf{100\scriptstyle{\pm 0}}$
ManiCM $\mathbf{100\scriptstyle{\pm 1}}$ $73\scriptstyle{\pm 3}$ $52\scriptstyle{\pm 3}$ $\mathbf{100\scriptstyle{\pm 1}}$ $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{93\scriptstyle{\pm 1}}$ $\mathbf{85\scriptstyle{\pm 11}}$ $\mathbf{100\scriptstyle{\pm 0}}$
Meta-World (Easy)
Method/Task Door Open Drawer Close Faucet Open Handle Pull Handle Press Side Handle Pull Side Lever Pull Plate Slide
DP3 [48] $86\scriptstyle{\pm 4}$ $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{100\scriptstyle{\pm 0}}$ $80\scriptstyle{\pm 7}$ $96\scriptstyle{\pm 0}$ $\mathbf{96\scriptstyle{\pm 4}}$ $\mathbf{90\scriptstyle{\pm 6}}$ $\mathbf{100\scriptstyle{\pm 0}}$
ManiCM $\mathbf{92\scriptstyle{\pm 1}}$ $\mathbf{100\scriptstyle{\pm 0}}$ $99\scriptstyle{\pm 2}$ $\mathbf{81\scriptstyle{\pm 7}}$ $\mathbf{97\scriptstyle{\pm 2}}$ $\mathbf{96\scriptstyle{\pm 3}}$ $88\scriptstyle{\pm 5}$ $\mathbf{100\scriptstyle{\pm 0}}$
Meta-World (Easy) Meta-World (Medium)
Method/Task Plate Slide Back Side Reach Reach Wall Window Close Peg Unplug Side Box Close Soccer Sweep Into
DP3 [48] $\mathbf{100\scriptstyle{\pm 0}}$ $68\scriptstyle{\pm 9}$ $\mathbf{80\scriptstyle{\pm 3}}$ $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{85\scriptstyle{\pm 16}}$ $79\scriptstyle{\pm 2}$ $\mathbf{22\scriptstyle{\pm 1}}$ $38\scriptstyle{\pm 19}$
ManiCM $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{77\scriptstyle{\pm 16}}$ $79\scriptstyle{\pm 5}$ $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{85\scriptstyle{\pm 17}}$ $\mathbf{81\scriptstyle{\pm 5}}$ $\mathbf{22\scriptstyle{\pm 1}}$ $\mathbf{40\scriptstyle{\pm 19}}$
Meta-World (Hard) Meta-World (Very Hard)
Method/Task Hand Insert Pick Out of Hole Pick Place Shelf Place Stick Pull Stick Push Pick Place Wall
DP3 [48] $\mathbf{25\scriptstyle{\pm 2}}$ $\mathbf{53\scriptstyle{\pm 5}}$ $\mathbf{56\scriptstyle{\pm 6}}$ $48\scriptstyle{\pm 6}$ $60\scriptstyle{\pm 3}$ $97\scriptstyle{\pm 5}$ $45\scriptstyle{\pm 3}$
ManiCM $\mathbf{25\scriptstyle{\pm 4}}$ $\mathbf{53\scriptstyle{\pm 3}}$ $55\scriptstyle{\pm 6}$ $\mathbf{55\scriptstyle{\pm 7}}$ $\mathbf{64\scriptstyle{\pm 2}}$ $\mathbf{98\scriptstyle{\pm 3}}$ $\mathbf{46\scriptstyle{\pm 6}}$
A.4 Societal Impacts
We discuss the potential societal impacts of our ManiCM in this section. For positive impacts, our open-sourced code can serve as a simple yet effective baseline to accelerate diffusion-based policies for the robotic community. Besides, in the pursuit of the ‘Embodied Foundation Models’, ManiCM can be directly extended to distill large-scale diffusion-based policies to real-time closed-loop control level. However, our method is essentially an end-to-end learning-based generative model. Thus unlike some methods like Task-and-Motion-Planning (TAMP) that exhibit strong theoretical guarantees on the expected risk, ManiCM may generate harmful actions when faced with purely unseen environments. Thus, we recommend mitigating the potential risk by setting a ‘safety region’ for the robot arm in real deployment. Our model learns the expert demonstrations from scratch, so it does not contain harmful biases embedded in the internet-scale pre-trained models.

### v2

@@ III-C Manipulation Consistency Distillation
To train ManiCM, we adopt the method of consistency distillation, as depicted in Figure 2. Here $\boldsymbol{a}_{0}$ represents the sequence of robotic arm actions, which comprises multiple actions, with the dimensionality determined by the specific task. $pcd$ denotes the point cloud obtained from images while $pose$ signifies the current pose of the robotic arm from the offline dataset. These inputs are processed through an encoder to obtain compact representations, which are then used as conditional inputs for the teacher network, online network, and target network. Given a ground-truth expert action $\boldsymbol{a}_{0}$, we first obtain the noisy action $\boldsymbol{a}_{n+k}$ by conducting a forward diffusion operation with $n+k$ steps to add noise to $\boldsymbol{a}_{0}$, where $k$ is the skipping interval mentioned in Section III-A. Subsequently, the noisy action $\boldsymbol{a}_{n+k}$ is feed-forwarded to a frozen teacher network and the online network to decode the clean $\hat{\boldsymbol{a}}_{0}^{*}$ and $\hat{\boldsymbol{a}}_{0}$, respectively. In order to make sure the self-consistency property Equation 1, we design a target network cloned from the online network, and we expect the denoised outputs of the online network and the target network to be aligned. Specifically, the target network generates $\hat{\boldsymbol{a}}_{0}^{-}$ utilizing $\boldsymbol{a}_{n}$ obtained from processing $\hat{\boldsymbol{a}}_{0}^{*}$ with a $k$-step ODE solver $\bf\Phi$ (e.g., DDIM [24]), such that $\hat{\boldsymbol{a}}_{0}$ and $\hat{\boldsymbol{a}}_{0}^{-}$ are obtained by the consistency function $\textbf{{f}}(\cdot,\cdot)$ in  Equation 2. Therefore, the consistency distillation loss is expressed as,
@@
where ${\bf\Phi}(\cdot,\cdot)$ is the update function of a $k$-step ODE solver applied to PF-ODE. As outlined in Section III-A, we utilize EMA to update the parameters of the target network from the parameters of the online network. We refer to the pre-trained manipulation 3D diffusion model, such as DP3 [6], as the "teacher network".
@@
Compared to other manipulation diffusion models, our ManiCM model achieves the fastest runtime ($\sim$16ms per action) during the inference phase by sampling high-quality actions in just one step, which is illustrated in Table I.
@@
IV Experiments
In this section, we first introduce the experiment setup including datasets, baseline methods, evaluation metrics and implementation details (Section IV-A). Then we compare our method with the state-of-the-art approaches to show the superiority in runtime efficiency and success rate on simulation environment (Section IV-B) and real-world environment(Section IV-C). Subsequently, we conduct an ablation study to explore different design choices in our ManiCM (Section IV-D). We also illustrate the visualization results to depict our intuition (Section IV-E). Further results can be found in the supplementary material.
IV-A Experimental Setup
Datasets. We conduct our single-task experiments in the well-recognized MetaWorld [44] and Adroit [45] benchmarks, resulting in a total of $31$ tasks. These tasks range from simple pick-and-place tasks to more challenging scenarios such as dexterous manipulation, which ensure that the model is effective across various scenarios. Specifically, MetaWorld tasks are grouped into different difficulty levels from easy to very hard according to [46]. For offline training datasets, we obtain expert demonstrations through heuristic policies in MetaWorld, and acquire them utilizing the reinforcement learning method VRL3 [47] in Adroit. We obtain a limited number of $10$ expert demonstrations for training in each benchmark, following [6] for fair comparisons. We also conduct multi-task experiments on eight high-quality tasks selected from RLBench [48] benchmark. All models are trained with 20 demonstrations per task and use only the front camera view following [29].
Table III: Ablation on Different Diffusion Steps of the Teacher Model. ‘SR’ refers to the success rate. $\rightarrow$’ means skipping from the original training denoising steps to $k$ steps in inference with DDIM-Solver.
@@ IV-A Experimental Setup
Baselines. This work primarily emphasizes accelerating the inference speed of 3D modality in diffusion policy to achieve real-time robotic action generation. To this end, our main baseline is the state-of-the-art point cloud-based 3D diffusion policy (DP3 [6]). Additionally, we compare our method with ‘Simple DP3’, a variation of DP3 with a lightweight policy backbone. Our baselines also include the well-known diffusion policy (DP [5]) and 3D Diffuser Actor (3dda [29]) for more comprehensive performance comparisons. To ensure fairness, the resolution of all images and depth maps remains consistent across all experiments.
Evaluation Metrics. Following [6], we assess $20$ episodes every $200$ training epochs and then calculate the average of the highest $5$ success rates, along with the average runtime per step. For each main experiment, we run $3$ random seeds with seed numbers 00, $1$, $2$ to mitigate the performance fluctuation. We also report the Number of Function Evaluations (NFE) as an additional device-agnostic metric for inference speed. For the success rate and runtime, the higher the better. For the NFE, the lower the better.
Implementation Details. For ManiCM, an AdamW [49] optimizer with a batch size of 128 and a learning rate of 5e-5 is employed for 3K epochs training. We also adopt a cosine decay learning rate scheduler and 500 iterations of linear warm-up. The EMA rate is set to $\mu=0.95$ by default and we use DDIM-Solver [24] with a skipping step of $k=10$ as ODE Solvers. Unless otherwise specified, our experiments are conducted using $100$ training time steps. The teacher model employed in our experiments uses DDIM with an inference step of $10$. For the input, We utilize an observation steps value of $2$, indicating that we use the point clouds from the two most recent time steps as conditions. Similarly, with a horizon of $4$, we predict four consecutive actions during each iteration for long-horizon tasks, while only $2$ actions are executed per prediction. To focus on the interest area, the image size is cropped to $84\times 84$ pixels. The point cloud is downsampled to $512$ points using the Farthest Point Sampling (FPS) algorithm, while the dimension of agent poses and actions are determined by the specific task. Both point clouds and agent poses are encoded using an MLP encoder into embeddings of length $64$. For normalization, we limit the min and max of action and observation to $\left[-1,1\right]$ independently since the prediction of DDIM would clip the prediction to $\left[-1,1\right]$ to ensure training stability. We implement our model in PyTorch [50], and all compared models are trained and evaluated on one NVIDIA RTX 4090 GPU to benchmark the decision runtime.
Figure 5: Multi-Task performance Comparison. ManiCM balances accuracy with 3D Diffuser Actor [29] while achieving ×13.3 faster inference speeds on RLBench.
IV-B Comparison with the State-of-the-Art Methods
Comparisons on Runtime. Table I illustrates the comparison of the runtime per step of each model. Our method achieves an impressive runtime around 16ms per step on average, which is $\sim$10$\times$ faster than the baseline method DP3. The acceleration comes from the manipulation consistency model, which can generate robot actions in only one-step inference. Besides, the $4$-step inference version of ManiCM also improves the efficiency by $\sim$3$\times$, Moreover, the proposed ManiCM also surpasses the lightweight version Simple DP3 significantly, which shows that the speedup from reducing the sampling speed is a more efficient way than shrinking the model itself. The results demonstrate the superiority of our ManiCM in terms of runtime, which verifies our intuition that incorporating the consistency model boosts the inference speed in various robotic manipulation tasks.
Comparisons on Success Rate. Table II depicts the comparison of the success of each model. The results show that our method accelerates the diffusion process without any performance drop. Specifically, with only one-step inference, our ManiCM can approximate or even surpass the state-of-the-art model DP3 and Simple DP3 by $1.3$%. By distilling the diffusion process with consistency regularization, our method avoids overfitting to the training demonstrations. Especially, in the Metaworld Very Hard tasks, our method outperforms the state-of-the-art method with an absolute improvement of $5.5$%. Moreover, increasing the sampling steps in ManiCM yields even better performance on average ($79.0$% vs. $78.5$%). The above experimental results prove the effectiveness of handling robotic tasks with our manipulation consistency model.
Comparisons on Multi-Task Setup. Figure 5 compares ManiCM with the 3D Diffuser Actor [29] baseline on the multi-task RLBbench simulation environment. The results show that ManiCM achieves competitive performance while demonstrating a remarkable 13.3× acceleration in inference speed, highlighting our consistency model’s effectiveness on a multi-task setup.
Figure 6: Our real-world experiment setup. We use an end effector and a DH Robotics AG-95 gripper based on UR3e arms and include three diverse manipulation tasks. A Realsense D435i camera is applied to capture visual observations.
Table IV: Real World Experiment Results. Each task is evaluated with 10 trials.
Method Slide Block Stack Cube Transfer Piece
SR Runtime SR Runtime SR Runtime
DP3 $\mathbf{90}$ $177.2$ $\mathbf{80}$ $181.3$ $70$ $172.4$
ManiCM $80$ $\mathbf{16.9}$ $\mathbf{80}$ $\mathbf{17.2}$ $\mathbf{80}$ $\mathbf{16.8}$
IV-C Real World Experiments
As shown in Figure 6, we evaluate ManiCM in the real world on three tasks from easy to hard with different embodiments: Slide Block, Stack Cube, and Transfer Piece.
  1. 1.
    Slide Block: The robot has to push the gray block off the white cube with the end-effector (see LABEL:subfig:slide_block).
  2. 2.
    Stack Cube: The robot has to pick up the red cube and then stack it on top of the white cube with a parallel jaw gripper (see LABEL:subfig:stack_cube).
  3. 3.
    Transfer Piece: The robot has to partially close the gripper, insert it into the left side of the box to pick the piece, and then place the piece on the right side of the box with a parallel jaw gripper. This task is more complex and requires higher precision (see LABEL:subfig:transfer_piece).
We use an end effector and a DH Robotics AG-95 gripper based on UR3e arms and a Realsense D435i camera to capture visual observations, with an NVIDIA RTX 4090 GPU for inference. We retain the same network architecture and input-output configuration as in the simulation environment. For each task, we conduct 10 trials to evaluate the average success rate and inference time. As demonstrated in Table IV, ManiCM achieves 10× faster inference speeds than DP3 while maintaining comparable success rates, validating its effectiveness in real-world setup.
IV-D Ablation Studies
@@
Ablation on Different Diffusion Steps of the Teacher Model. We compare our ManiCM with the state-of-the-art method DP3 under different DDIM sampling steps in Table III. We can conclude that: 1) The task success rates do not monotonically increase when taking more diffusion sampling steps in both DP3 and ManiCM, which indicates that most of the diffusion steps are redundant. As a result, we can distill the teacher model with multiple-step inference into an online model that generates high-quality action in only one sampling step. 2) For different diffusion sampling steps, our ManiCM can stably distill competitive capacity in task completion, where it accelerates the runtime efficiency by $\sim$57 $\times$ when the teacher is a DDIM with $50$ inference steps.
IV-E Qualitative Analysis
@@
V Conclusion
In this paper, we have proposed a ManiCM method that imposes the consistency constraint on the diffusion process, which brings 3D diffusion-based robotic manipulation to a real-time level. We design a manipulation consistency self-consistency function that predicts the action sample directly based on 3D point cloud condition. Subsequently, we implement a manipulation consistency distillation to ensure the action is denoised from any point along the ODE trajectory within only one-step inference, which avoids the inefficient iterative denoising procedure in the original diffusion model. Extensive experiments across various task suites verify the effectiveness and efficiency of our method. However, the scalability of the consistency model to distill a large-scale pre-trained diffusion policy is still under-explored and we leave this issue for future work.